# On Mirzakhani’s twist torus conjecture

### v2

@@ 1. Introduction
Twist tori lift to $\mathcal{P}^{1}\mathcal{M}_{g}$ by keeping track of the multicurve $\gamma$ weighted by $(L\cdot\#\gamma)^{-1}$. Evidently, these lifted twist tori are invariant under the earthquake flow. Conjecture 1.1 would therefore follow from the stronger statement that the normalized Lebesgue measures on lifted twist tori equidistribute to $\mu_{\operatorname{Mirz}}/b_{g}$ in $\mathcal{P}^{1}\mathcal{M}_{g}$ as $L\to\infty$.
@@
Say that a pair of pants has property ()($\nabla$)( ∇ ) if one of its boundary lengths is the sum of the other two.
@@
Fix any pants decomposition $\gamma$, any >$\bm{\ell}\in\mathbb{R}^{3g-3}_{>0}$, and any $\mathbf{h}\in\mathbb{R}^{3g-3}_{>0}$ such that $\bm{\ell}\cdot\mathbf{h}=1$. Then there is a set $Z\subset\mathbb{R}$ of zero density such that
$\lim_{\begin{subarray}{c}t\to\infty\\ t\notin Z\end{subarray}}\mu_{\gamma}(e^{t}\bm{\ell},e^{-t}\mathbf{h})=\mu_{\infty}$
@@
    if $g\geq 3$ and no component of $S\setminus\gamma$ has ()($\nabla$)( ∇ ), then $\mu_{\infty}$ is $\mu_{\operatorname{Mirz}}/b_{g}$.
@@
    if some component of $S\setminus\gamma$ has ()($\nabla$)( ∇ ), then $\mu_{\infty}$ is not equal (hence singular) to $\mu_{\operatorname{Mirz}}/b_{g}$.
In particular, Conjecture 1.1 is true outside a set of times $e^{Z}$ where $Z$ has zero density.
Remark 1.3.
For $g=2$, statement (1) of Theorem 1.2 would follow, using our techniques, from [AW23, Conjecture 1.2] about “almost high rank” invariant submanifolds; see §§45.
@@
The limiting measure $\mu_{\infty}$ always has more structure than just earthquake-flow invariance. In [CF24b], the authors extended the earthquake flow to a (measurable) action by “stretchquakes.” There and in [CF24a], the authors showed that stretchquake-invariant probabilty measures on $\mathcal{P}^{1}\mathcal{M}_{g}$ correspond to Lebesgue-class measures on affine $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$-invariant subvarieties of strata of holomorphic quadratic differentials (§4.2). The measure $\mu_{\infty}$ is always stretchquake-invariant, and most of the time, we can identify it as corresponding to the Masur–Smillie–Veech measure on a component of a stratum of quadratic differentials with some zeros of order 2; see Corollary 5.12. We emphasize that Mirzakhani’s conjecture only references $\mu_{\operatorname{Mirz}}$; it seems to have been unexpected that families of twist tori could equidistribute to anything else.
@@
The strategy of our proof is to exploit a deep connection between the earthquake and horocycle flows discovered by Mirzakhani [Mir08] and further developed by the authors in [CF24b, CF24a]. Briefly, there is a Borel isomorphism $\mathcal{O}:\mathcal{P}^{1}\mathcal{M}_{g}\to\mathcal{Q}^{1}\mathcal{M}_{g}$ that takes the earthquake flow to the horocycle flow on the moduli space $\mathcal{Q}^{1}\mathcal{M}_{g}$ of unit area quadratic differentials in a time-preserving way. While $\mathcal{O}$ is not continuous, it is continuous along the leaves of a certain foliation and each $\mathbb{T}_{\gamma}(\bm{\ell},\mathbf{h})$ is contained in such a leaf. This allows us to map the twist torus via $\mathcal{O}$ to get a torus in $\mathcal{Q}^{1}\mathcal{M}_{g}$ consisting of flat surfaces glued together from horizontal cylinders corresponding to the components of $\gamma$ with heights $\mathbf{h}$ and lengths $\bm{\ell}$; see Figure 1. Once in this setting, we can use powerful results from Teichmüller dynamics to understand the limiting distribution of expanding twist tori in $\mathcal{Q}^{1}\mathcal{M}_{g}$.
@@
Thus, having proved an equidistribution statement for expanding twist tori in $\mathcal{Q}^{1}\mathcal{M}_{g}$, we can simply pull it back to deduce Theorem 1.2. This scheme also works to show the equidistribution of many different families of expanding twist tori where $\gamma$ is an arbitrary multicurve, and can be used to show equidistribution to many different limiting distributions. See Theorems 5.13, 5.14, and 5.16.
@@ 1.1. Context: unipotent flows on moduli spaces
For a lattice $\Gamma\leq G$ and horospherical subgroup $U\leq G$, the $U$-invariant ergodic probability measures and $U$-orbit closures in $\Gamma\backslash G$ were classified by Dani [Dan86]: they are homogeneous. In her landmark results, Ratner proved that such rigidity phenomena hold for arbitrary connected subgroups $H\leq G$ generated by unipotents [Rat91a, Rat91b]: every $H$-invariant ergodic probability measure and $H$-orbit closure in $\Gamma\backslash G$ is homogeneous.
@@
Leveraging the ergodicity of the earthquake flow, Mirzakhani showed that the analogues of expanding horospheres ($L$-level sets of the hyperbolic length of a simple closed curve) equidistribute to $\pi_{*}\mu_{\operatorname{Mirz}}$ on $\mathcal{M}_{g}$ as $L\rightarrow\infty$ [Mir07]. These level sets can be lifted to $\mathcal{P}^{1}\mathcal{M}_{g}$ in a natural way by keeping track of the curve, as in §3.1. There are finite, earthquake flow–invariant measures on these lifted level sets that are uniformly expanded/contracted in forward/backward time by the “dilation flow,” the diagonal part of the stretchquake action (see the end of §2.3). The lifted level sets have the maximum dimension with this property; see also Theorems 2.1 and 2.3 that explain that the conjugacy $\mathcal{O}$ maps lifted level sets to leaves of the strong unstable foliation for the Teichmüller geodesic flow. These observations justify Mirzakhani’s use of the “horospherical” terminology in [Mir07] (compare [Wri20, §13.3]).
@@
Twist tori are only $(3g-3)$-dimensional inside of the $(6g-7)$-dimensional expanding horospheres of the previous paragraph; as such, they should not be considered horospherical. The much more refined results of Eskin–Mirzkhani–Mohammadi, analogous to Ratner’s Theorems, are vital ingredients in our proof of Theorem 1.2 and the even more delicate examples presented in §5.4. Compare Figure 2.
@@
Mirzakhani’s equidistribution results for expanding horospheres were generalized by both Arana-Herrera [AH21] (who clarified a technical point in [Mir07]) and Liu [Liu20] to different types of length functions associated to multicurves and by Arana-Herrera and the first author [AHC22] to the case where one puts restrictions on the geometry of the complementary subsurface.
@@
Our techniques recover and generalize these results. Indeed, let $\gamma=\gamma_{1}\cup\ldots\cup\gamma_{k}$ be any multicurve and $\mathbf{h}\in\mathbb{R}^{k}_{>0}$ any weight vector. Consider the $L$-level set of the hyperbolic length function for $\mathbf{h}\gamma:=h_{1}\gamma_{1}\cup\ldots\cup h_{k}\gamma_{k}$ in the Teichmüller space $\mathcal{T}_{g}$; this admits a lift to the unit cotangent bundle $\mathcal{P}^{1}\mathcal{T}_{g}$ by setting the second coordinate equal to $\mathbf{h}\gamma/L$. Any probability measure supported on such a lifted level set that is absolutely continuous with respect to the Lebesgue class with continuous density and such that its restriction to each earthquake flow line is the restriction of Lebesgue to an interval is said to be horospherical. These level sets can also be stratified by the combinatorial type of the ribbon graph spine for $X\setminus\gamma$ (see §2 and §§3.23.3); any measure satisfying the properties above but on a substratum is called a horospherical slice measure. We use the same terms to denote the pushforwards of such measures down to $\mathcal{P}^{1}\mathcal{M}_{g}$. See Section 3.3 and Figure 2.
@@
In the second case, the limiting measure is the pushforward via $\mathcal{O}^{-1}$ of the Masur–Smillie–Veech measure on the stratum component $\mathcal{Q}$ of $\mathcal{Q}^{1}\mathcal{M}_{g}$ corresponding to the horospherical slice; see §4.3 for the proof. The generic pair $(X,\lambda)$ with respect to one of these measures is a hyperbolic surface equipped with a measured lamination that cuts it into a union of regular ideal polygons [CF24b, Corollary 1.2].
@@ Acknowledgments
We are very grateful to Francisco Arana-Herrera for an inspiring discussion at the beginning of this project and to Paul Apisa for explaining to us how to deduce Corollary 5.4 from his work with Alex Wright. AC was supported by NSF grants DMS-2005328 and DMS-2202703. JF was supported by NSF grant DMS-2005328 and DFG grants 427903332 and 281071066 – TRR 191.
@@ 2. Flows on moduli spaces
In this section, we discuss basic dynamical features of the spaces $\mathcal{P}^{1}\mathcal{M}_{g}$ and $\mathcal{Q}^{1}\mathcal{M}_{g}$ and describe a Borel measurable bijection $\mathcal{O}:\mathcal{PM}_{g}\to\mathcal{QM}_{g}$ conjugating their dynamics. Throughout, $S$ is a closed, oriented surface of genus $g\geq 2$ and $\mathcal{T}_{g}$ is the Teichmüller space of marked hyperbolic or conformal structures on $S$. The mapping class group of $S$ will be denoted $\operatorname{Mod}_{g}$, and $\mathcal{M}_{g}=\mathcal{T}_{g}/\operatorname{Mod}_{g}$ is its moduli space.
@@ 2.1. The horocycle flow
Given a Riemann surface $X\in\mathcal{T}_{g}$, complex deformation theory identifies the cotangent space $T_{X}^{*}\mathcal{T}_{g}$ with the vector space of holomorphic quadratic differentials $Q(X)$ on $X$. Let $\mathcal{QT}_{g}$ denote the bundle of non-zero quadratic differentials over $\mathcal{T}_{g}$ and $\mathcal{QM}_{g}=\mathcal{QT}_{g}/\operatorname{Mod}_{g}$. By integrating ±$\pm\sqrt{q}$, each quadratic differential $q$ defines a flat metric on $q\setminus Z(q)$, where $Z(q)$ is the set of zeros of $q$. The area is given by $\int|q|$, and we denote by $\mathcal{Q}^{1}\mathcal{T}_{g}$ and $\mathcal{Q}^{1}\mathcal{M}_{g}$ the unit area loci, which can be identified with the unit cotangent bundles of $\mathcal{T}_{g}$ and $\mathcal{M}_{g}$, respectively.
Both $\mathcal{QT}_{g}$ and $\mathcal{QM}_{g}$ and their unit area loci are stratified by the number and orders of zeros of the differentials. Every partition $\underline{\kappa}=(\kappa_{1},...,\kappa_{n})$ of $44g-$ 4 and $\varepsilon\in\{\pm 1\}$ determines a (possibly empty) stratum $\mathcal{QT}_{g}(\underline{\kappa};\varepsilon)$ of quadratic differentials that are ($\varepsilon=+1$) or are not ($\varepsilon=-1$) squares of abelian differentials. Equivalently, the completion of the flat metric defined by $q\in\mathcal{QT}_{g}(\underline{\kappa};\varepsilon)$ has $n$ cone points, each of which has cone angle excess $\kappa_{i}\pi$. The rotational holonomy of the metric lives in $\mathbb{Z}/2\mathbb{Z}$ and is trivial if and only if $\varepsilon=+1$. We use $\mathcal{QM}_{g}(\underline{\kappa};\varepsilon)$ to denote a stratum of $\mathcal{QM}_{g}$ and add a $1$ superscript to denote unit area loci. Strata are not necessarily connected, but at least at the level of moduli space their connected components are classified [KZ03, Lan08, CM14].
Strata are complex orbifolds of dimension $(2g-2)+n+\delta$, where $\delta=1$ if $\varepsilon=+1$ and $\delta=0$ otherwise. This can be seen via local period coordinates, constructed as follows. Every $q$ has a a holonomy double cover $\widehat{q}\to q$, branched over the zeros of odd order, so that $\widehat{q}$ is a (possibly disconnected) abelian differential with a holomorphic covering involution. Let $H^{1}(\widehat{S},\widehat{Z};\mathbb{C})^{-}$ denote the $(-1)$-subspace for this involution, where $(\widehat{S},\widehat{Z})\cong(\widehat{q},Z(\widehat{q}))$. Take a triangulation of $q$ by saddle connections, singularity-free geodesics joining the cone points of $q$; this can be lifted to a triangulation $\widehat{\mathsf{T}}$ of $\widehat{q}$ by saddle connections. Integrating $\widehat{q}$ over the edges of $\widehat{\mathsf{T}}$ defines a cohomology class in $H^{1}(\widehat{S},\widehat{Z};\mathbb{C})^{-}$, and any small deformation of all of the triangles subject to the constraints imposed by $H^{1}(\widehat{S},\widehat{Z};\mathbb{C})^{-}$ defines a new quadratic differential in the same stratum. The period vector mapping into $H^{1}(\widehat{S},\widehat{Z};\mathbb{C})^{-}$ near $q$ is called a period coordinate chart. Two overlapping period coordinate charts differ by an integer linear transformation. We refer the reader to [Zor06], [Wri20], and [AM24] for a more thorough discussion of period coordinates and background on flat surfaces.
@@
$T_{q}\mathcal{QT}_{g}(\underline{\kappa};\varepsilon)\cong H^{1}(\widehat{S},\widehat{Z};\mathbb{C})^{-}.$
@@
There is an action of $\mathsf{SL}_{2}\mathbb{R}$ on $\mathcal{Q}^{1}\mathcal{M}_{g}$ preserving each stratum component $\mathcal{Q}$. For $M\in\mathsf{SL}_{2}\mathbb{R}$ and $q\in\mathcal{Q}$, $M\cdot q\in\mathcal{Q}$ is defined by postcomposing charts for the singular flat metric obtained by integration with the $\mathbb{R}$-linear map $M:\mathbb{C}\to\mathbb{C}$. Since these charts are only defined up to ±$\pm 1$, this $\mathsf{SL}_{2}\mathbb{R}$ action factors through $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$. In a period coordinate chart $H^{1}(\widehat{S},\widehat{Z};\mathbb{C})^{-}\cong\mathbb{C}\otimes H^{1}(\widehat{S},\widehat{Z};\mathbb{R})^{-}$, the action is given by applying $M$ to $\mathbb{C}$. The one-parameter subgroups
@@
On each component $\mathcal{Q}$ of each stratum of quadratic differentials, there is an $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$-invariant ergodic Masur–Smillie–Veech probability measure $\nu_{\mathcal{Q}}$ in the class of Lebesgue. It is constructed from the natural affine measure in period coordinate charts by assigning to a set $E\subset\mathcal{Q}$ the affine measure of the cone $\{tE\mid t\in(0,1]\}$. A characterization of the $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$-invariant ergodic probability measures on $\mathcal{Q}^{1}\mathcal{M}_{g}$ was given by Eskin–Mirzakhani [EM18]. See §4.2 and Theorem 4.2 below.
@@
Following [Vee86] and [For02] (see also [ABEM12, Theorem 3.15]), we define the strong stable and strong unstable foliations for the Teichmüller geodesic flow as follows. At every point $q$ the strong stable (unstable) leaf through $q$ locally consists of those $q^{\prime}$ with the same real (imaginary) foliation as $q$. Each unstable leaf
$\mathcal{W}^{uu}(\lambda):=\{q\in\mathcal{QT}_{g}:\text{ the imaginary foliation }\operatorname{Im}(q)\text{ is measure equivalent to }\lambda\}$
therefore has a stratified real-analytic structure coming from restricting the linear structure of strata of $\mathcal{QT}_{g}$ to intersections with $\mathcal{W}^{uu}(\lambda)$. For every $\lambda$, define $\mathcal{MF}(\lambda):=\{\eta\in\mathcal{MF}_{g}\mid(\eta,\lambda)\notin\Delta\}.$ Using (1), we have

### v3

@@ 1. Introduction
Twist tori lift to $\mathcal{P}^{1}\mathcal{M}_{g}$ by keeping track of the pants decomposition $\gamma$ weighted by 1/(3g3)L1/(3g-3)L. Evidently, these lifted twist tori are invariant under the earthquake flow. Conjecture 1.1 would therefore follow from the stronger statement that the normalized Lebesgue measures on lifted twist tori equidistribute to $\mu_{\operatorname{Mirz}}/b_{g}$ in $\mathcal{P}^{1}\mathcal{M}_{g}$ as $L\to\infty$.
@@
Say that a pair of pants has property $(\nabla)$ if one of its boundary lengths is the sum of the other two. Pick $(X,\mathbf{h}\gamma)\in\mathcal{P}\mathbb{T}_{\gamma}(\bm{\ell},\mathbf{h})$, and let $\nabla_{\gamma}(\bm{\ell})$ denote the number of components of $X\setminus\gamma$ that have $(\nabla)$; this number does not depend on the choice of $X$ or $\mathbf{h}$.
@@
Fix any pants decomposition $\gamma$, any $\bm{\ell}\in\mathbb{R}^{3g-3}_{>0}$, and any $\mathbf{h}\in\mathbb{R}^{3g-3}_{>0}$ such that $\bm{\ell}\cdot\mathbf{h}=1$. There is an earthquake flow invariant probability measure $\mu_{\infty}$ on $\mathcal{P}^{1}\mathcal{M}_{g}$ and a set $Z\subset\mathbb{R}$ of zero density such that
$\lim_{\begin{subarray}{c}t\to\infty\\ t\notin Z\end{subarray}}\mu_{\gamma}(e^{t}\bm{\ell},e^{-t}\mathbf{h})=\mu_{\infty}.$
The measure $\mu_{\infty}$ arises from the Masur–Smillie–Veech measure on a component of a stratum of quadratic differentials having $4g-4-2\nabla_{\gamma}(\bm{\ell})$ simple zeros and $\nabla_{\gamma}(\bm{\ell})$ zeros of order $2$. In particular,
@@
    if $\nabla_{\gamma}(\bm{\ell})=0$, then $\mu_{\infty}$ is $\mu_{\operatorname{Mirz}}/b_{g}$.
@@
    if $\nabla_{\gamma}(\bm{\ell})>0$, then $\mu_{\infty}$ is not equal (hence singular) to $\mu_{\operatorname{Mirz}}/b_{g}$.
Consequently, Conjecture 1.1 is true outside a set of times $e^{Z}$ where $Z$ has zero density.
We emphasize that Mirzakhani’s conjecture only references $\mu_{\operatorname{Mirz}}$; it seems to have been unexpected that families of twist tori could equidistribute to anything else. See Theorem 5.6 for a detailed, precise formulation of the limiting distributions that appear in Theorem 1.2. The generic pair $(X,\lambda)$ with respect to one of these measures is a hyperbolic surface equipped with a measured lamination that cuts it into a union of regular ideal polygons [CF24b, Corollary 1.2].
@@
$\frac{1}{T}\int_{0}^{T}\frac{\tau_{\gamma}(e^{t})}{e^{(3g-3)t}}~dt\to\frac{B(X)}{b_{g}}d\operatorname{vol}_{\operatorname{WP}}(X).$
@@
Thus, having proved an equidistribution statement for expanding twist tori in $\mathcal{Q}^{1}\mathcal{M}_{g}$, we can simply pull it back to deduce Theorem 1.2. This scheme also works to show the equidistribution of many different families of expanding twist tori where $\gamma$ is an arbitrary multicurve, and can be used to show equidistribution to many different limiting distributions. See Section 6, in particular Theorems 6.2, 6.9, and 6.10, the last of which shows that $2g$-dimensional expanding twist tori often equidistribute to Mirzakhani measure.
Remark 1.4.
One could also ask about expanding twist tori in $\mathcal{M}_{g,n}$, that is, when the hyperbolic surfaces are allowed to have cusps. Our methods rely on all of our previous results on the conjugacy $\mathcal{O}$, which do not yet have analogues in the punctured setting.
@@ 1.1. Context: unipotent flows on moduli spaces
For a lattice $\Gamma\leq G$ and horospherical subgroup $U\leq G$, the $U$-invariant ergodic probability measures and $U$-orbit closures in $\G\Gamma\backslash$ G were classified by Dani [Dan86]: they are homogeneous. In her landmark results, Ratner proved that such rigidity phenomena hold for arbitrary connected subgroups $H\leq G$ generated by unipotents [Rat91a, Rat91b]: every $H$-invariant ergodic probability measure and $H$-orbit closure in $\G\Gamma\backslash$ G is homogeneous.
@@
Leveraging the ergodicity of the earthquake flow, Mirzakhani showed that the analogues of expanding horospheres ($L$-level sets of the hyperbolic length of a simple closed curve) equidistribute to $\pi_{*}\mu_{\operatorname{Mirz}}$ on $\mathcal{M}_{g}$ as $L\rightarrow\infty$ [Mir07]. These level sets can be lifted to $\mathcal{P}^{1}\mathcal{M}_{g}$ in a natural way by keeping track of the curve, as in §3.1. There are finite, earthquake flow–invariant measures on these lifted level sets that are uniformly expanded/contracted in forward/backward time by the “dilation flow,” the diagonal part of the stretchquake action (see the end of §2.3). The lifted level sets have the maximum dimension with this property; see also Theorems 2.1 and 2.3 that explain that the conjugacy $\mathcal{O}$ maps lifted level sets to leaves of the unstable foliation for the Teichmüller geodesic flow. These observations justify Mirzakhani’s use of the “horospherical” terminology in [Mir07] (compare [Wri20, §13.3]).
@@
Twist tori are only (3g3)(3g-3)-dimensional inside of the (6g7)(6g-7)-dimensional expanding horospheres of the previous paragraph; as such, they should not be considered horospherical. The much more refined results of Eskin–Mirzkhani–Mohammadi, analogous to Ratner’s Theorems, are vital ingredients in our proof of Theorem 1.2 and the even more delicate examples presented in §6. Compare Figure 2.
@@
Mirzakhani’s equidistribution results for expanding horospheres were generalized by both Arana-Herrera [AH21] (who clarified a technical point in [Mir07]) and Liu [Liu20] to different types of length functions associated to multicurves and by Arana-Herrera and the first author [AHC22, Theorem 5.2] to the case where one puts restrictions on the geometry of the complementary subsurface.
@@
Our techniques recover and generalize these equidistribution results. Indeed, let $\gamma=\gamma_{1}\cup\ldots\cup\gamma_{k}$ be any multicurve and $\mathbf{h}\in\mathbb{R}^{k}_{>0}$ any weight vector. Consider the $L$-level set of the hyperbolic length function for $\mathbf{h}\gamma:=h_{1}\gamma_{1}\cup\ldots\cup h_{k}\gamma_{k}$ in the Teichmüller space $\mathcal{T}_{g}$. A measure on this level set is horospherical if it is absolutely continuous with respect to the Lebesgue class with continuous density, and if the restriction to each earthquake flow line is Lebesgue measure. These level sets can also be stratified by the combinatorial type of the ribbon graph spine for $X\setminus\gamma$ (see §2 and §§3.23.3). Fixing a combinatorial type defines a submanifold of the $L$-level set, and any measure which is both absolutely continuous with respect to Lebesgue on such a submanifold and Lebesgue on earthquake flow orbits is called a horospherical slice measure. See Section 3.3 and Figure 2.
Both horospherical and horospherical slice measures admit earthquake-flow invariant lifts to the unit cotangent bundle $\mathcal{P}^{1}\mathcal{T}_{g}$ by setting the second coordinate equal to $\mathbf{h}\gamma/L$. These can further be pushed forward to $\mathcal{P}^{1}\mathcal{M}_{g}$, and we continue to use the same terminology.
@@
In the second case, the limiting measure arises via the Masur–Smillie–Veech measure on the component of the stratum of $\mathcal{Q}^{1}\mathcal{M}_{g}$ corresponding to the horospherical slice; see §4.3 for the proof.
@@ Acknowledgments
We are very grateful to Francisco Arana-Herrera for an inspiring discussion at the beginning of this project and to Paul Apisa for explaining his work with Alex Wright to us. We heartily thank the referee for supplying a more direct proof of the identification of orbit closures in the setting of pants decompositions. AC was supported by NSF grants DMS-2005328 and DMS-2202703. JF was supported by NSF grant DMS-2005328 and DFG grants 427903332 and 281071066 – TRR 191.
@@ 2. Flows on moduli spaces
In this section, we discuss basic dynamical features of the spaces $\mathcal{P}^{1}\mathcal{M}_{g}$ and $\mathcal{Q}^{1}\mathcal{M}_{g}$ and describe a Borel measurable bijection $\mathcal{O}$ conjugating their dynamics. Throughout, $S$ is a closed, oriented surface of genus $g\geq 2$ and $\mathcal{T}_{g}$ is the Teichmüller space of marked hyperbolic or conformal structures on $S$. The mapping class group of $S$ will be denoted $\operatorname{Mod}_{g}$, and $\mathcal{M}_{g}=\mathcal{T}_{g}/\operatorname{Mod}_{g}$ is the moduli space.
@@ 2.1. The horocycle flow
Given a Riemann surface $X\in\mathcal{T}_{g}$, complex deformation theory identifies the cotangent space $T_{X}^{*}\mathcal{T}_{g}$ with the vector space of holomorphic quadratic differentials $Q(X)$ on $X$. Let $\mathcal{QT}_{g}$ denote the bundle of non-zero quadratic differentials over $\mathcal{T}_{g}$ and $\mathcal{QM}_{g}=\mathcal{QT}_{g}/\operatorname{Mod}_{g}$. By integrating $\pm\sqrt{q}$, each quadratic differential $q$ defines a flat metric on $q\setminus Z(q)$, where $Z(q)$ is the set of zeros of $q$. The area is given by $\int|q|$, and we denote by $\mathcal{Q}^{1}\mathcal{T}_{g}$ and $\mathcal{Q}^{1}\mathcal{M}_{g}$ the unit area loci, which can be identified with the unit cotangent bundles of $\mathcal{T}_{g}$ and $\mathcal{M}_{g}$, respectively.
Both $\mathcal{QT}_{g}$ and $\mathcal{QM}_{g}$ and their unit area loci are stratified by the number and orders of zeros of the differentials. Every partition $\underline{\kappa}=(\kappa_{1},...,\kappa_{n})$ of 4g44g-4 and $\varepsilon\in\{\pm 1\}$ determines a (possibly empty) stratum $\mathcal{QT}_{g}(\underline{\kappa};\varepsilon)$ of quadratic differentials that are ($\varepsilon=+1$) or are not ($\varepsilon=-1$) squares of abelian differentials. Equivalently, the completion of the flat metric defined by $q\in\mathcal{QT}_{g}(\underline{\kappa};\varepsilon)$ has $n$ cone points, each of which has cone angle excess $\kappa_{i}\pi$. The rotational holonomy of the metric lives in $\mathbb{Z}/2\mathbb{Z}$ and is trivial if and only if $\varepsilon=+1$. We use $\mathcal{QM}_{g}(\underline{\kappa};\varepsilon)$ to denote a stratum of $\mathcal{QM}_{g}$ and add a $1$ superscript to denote unit area loci. Strata are not necessarily connected, but at least at the level of moduli space their connected components are classified [KZ03, Lan08, CM14].
[…]
@@
[…]
@@
[…]
@@
Each component $\mathcal{Q}$ of each stratum of quadratic differentials has a natural affine measure coming from period coordinate charts. This induces a $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$-invariant ergodic Masur–Smillie–Veech probability measure $\nu_{\mathcal{Q}}^{1}$ on the unit area locus $\mathcal{Q}^{1}$ by assigning to a set $E\subset\mathcal{Q}^{1}$ the affine measure of the cone $\{tE\mid t\in(0,1]\}$, divided by the total mass of $\mathcal{Q}^{1}$. A characterization of the $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$-invariant ergodic probability measures on $\mathcal{Q}^{1}\mathcal{M}_{g}$ was given by Eskin–Mirzakhani [EM18]. See §4.2 and Theorem 4.4 below.
@@
Following [Vee86] and [For02] (see also [ABEM12, Theorem 3.15]), we define a pair of transverse foliations of $\mathcal{QT}_{g}$ and $\mathcal{Q}^{1}\mathcal{T}_{g}$. Given $q\in\mathcal{QM}_{g}$, define
[…]
The intersection of $\mathcal{W}^{uu}(\lambda)$ with the unit area locus is often referred to as the strong unstable foliation for the Teichmüller geodesic flow. The foliation $\mathcal{W}^{ss}(\lambda)$ is defined similarly with the imaginary part replaced with the real, and its intersection with $\mathcal{Q}^{1}\mathcal{T}_{g}$ is often referred to as the strong stable foliation. Each leaf $\mathcal{W}^{uu}(\lambda)$ has a stratified real-analytic structure coming from restricting the linear structure of strata of $\mathcal{QT}_{g}$ to intersections with $\mathcal{W}^{uu}(\lambda)$.
For every $\lambda$, define $\mathcal{MF}(\lambda):=\{\eta\in\mathcal{MF}_{g}\mid(\eta,\lambda)\notin\Delta\}.$ Using (1), we have